The sidewall of a cylindrical tank of radius 30 cm and height of 1 m should be lined with a 4 cm thick layer of concrete.


The sidewall of a cylindrical tank of radius 30 cm and height of 1 m should be lined with a 4 cm thick layer of concrete.


Calculate the volume of the cylinder

Answer: What is approximately the amount of concrete required?


1
Expert's answer
2022-02-27T12:05:33-0500

Volume of the tank: "V_1=\pi r^2h=\pi*0.3^2*1=0.09\pi \;m^3."

Volume of the tank with concrete: "V_2=\pi (r+0.04)^2h=\pi*0.34^2*1=0.1156\pi\;m^3" .

Volume of the concrete: "V=V_2-V_1=0.1156\pi-0.09\pi=0.0256\pi\;m^3\approx 0.08\;m^3."